1) Four men A, B, C and D want to cross a bridge. They all begin on the same side. It is night, and they have only one flashlight with them. At most two men can cross the bridge at a time, and any party who crosses, either one or two people, must have the flashlight with them. The flashlight must be walked back and forth: it cannot be thrown, etc. Each man walks at a different speed. A pair must walk together at the speed of the slower man. A needs 1 minute to cross the bridge. B needs 2 minutes. C needs 5 minutes and D needs 10 minutes. For example, if man A and man C walk across together, they need 5 minutes. How can all four men cross the bridge in 17 minutes?
Answer #1

According to the given conditions, the solution to the problem is-

  1. A and B cross the bridge and A returns back with the flashlight: time taken= 2+1 = 3 min; Total time = 0+3 = 3 min
  2. C and D cross the bridge and B returns back with the flashlight: time taken= 10+2 = 12 min; Total time = 3+12 = 15 min
  3. Finally, A and B cross the bridge: time taken= 2 min; Total time = 15+2 = 17 min

Therefore the total time taken by this process is 17 min and all the 4 men are on the other side.